THE STRUCTURE OF THE 1-CHLORO-1-FLUOROETHYLENE–ACETYLENE COMPLEX

HELEN O. LEUNG, \underline{MARK D. MARSHALL}, AND DAVID D. GRIMES, Department of Chemistry, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000.

The insensitivity of the moments of inertia to certain geometric parameters for the limited set of isotopologues studied led to unexpected results for a preliminary structure of the complex formed between 1-chloro-1-fluoroethylene and acetylene presented three years ago.^{a} The availability of this chlorofluoroethylene with a single deuterium substitution in both (E) and (Z ) isomeric forms^{b} allows us to supplement the original data set with microwave spectra of (E)-CHDC^{35}ClF– HCCH and (Z )-CHDC^{35}ClF–HCCH. Supported by the results of ab initio calculations, a planar, chemically reasonable structure is obtained for this complex, in which the acetylene forms a hydrogen bond with the fluorine atom of 1-chloro-1-fluoroethylene. The hydrogen bond bends to allow a secondary interaction with the hydrogen atom cis to this fluorine atom. Comparisons with the structure of the analogous complexes formed with vinyl fluoride and 1,1-difluoroethylene reveal the effects of chlorine substitution geminal to the acceptor atom of the hydrogen bond.

aH.O. Leung and M.D. Marshall, The 62nd OSU International Symposium on Molecular Spectroscopy, Talk RG10 (2007)

bKindly provided by Prof. Norman Craig, Oberlin College